Consider the following program specification for evaluating a polynomial anxh + an- xn-1 + ... + a1x + ao at x = c: Input: c,(5) iti+1 (6) return y Implementation 2: Polynomial_2 Polynomial_2(c, ao, a1, ... , an) (1) it-1 (2) ytan (3) whileisn yty* cQuestion 1 4 pts Consider evaluating the polynomial x4 + 3x3 + 2x2 + 2x+1 at x = 3. Trace the execution of Polynomial_1 on thPlease answer the 2 blanks in Q1, show full working!! I want to know how this works, thanks. This is a discrete math question.
Answer #1

For polynomial_1

The first time it is executed is when i=1, y=a(0)=1 (constant term) so that y={\color{Red} 3}

The second time it is executed is when y + y +a; * Power(c, i) = 3+2* Power(3,1)

Which is y = 3+2+31 = 3+2:3=9

This algorithm is called Horner's method for polynomial evaluation

\blacksquare

Hope this was helpful. Please do leave a positive rating if you liked this answer. Thanks and have a good day!